Show that the function f: R^2→R^2 given by

f(x,y) = (xy^3+1, x^2+y^2) is not invertible. Futher check whether it is locally invertible at the point (2,1)


1
Expert's answer
2021-06-24T18:13:00-0400

Given function f: R2 "\to" R2

such that f(x,y) = (xy3+1, x2+y2)

we have to show that f is not invertible in R2


Let u = xy3+1,v = x2+y2


Df(x,y) = "\begin{vmatrix}
 \frac{\delta u}{\delta x} & \frac{\delta u}{\delta y} \\
 \frac{\delta v}{\delta x} & \frac{\delta v}{\delta y}
\end{vmatrix}"


Df(x,y) = "\begin{vmatrix}
 \frac{\delta }{\delta x}(xy^3+1) & \frac{\delta }{\delta y}(xy^3+1) \\
 \frac{\delta }{\delta x}(x^2+y^2) & \frac{\delta }{\delta y}(x^2+y^2)
\end{vmatrix}"


Df(x,y) = "\begin{vmatrix}
 y^3 & 3xy^2 \\
 2x & 2y
\end{vmatrix}"


Df(x,y) = 2y4"-" 6x2y2

Df(x,y) = 2y2(y2"-" 3x2)

here we see that, Df(x,y) = 0 if 2y2(y2"-" 3x2) = 0 "\implies" either y = 0 or (y2"-" 3x2) = 0 "\implies" either y = 0 or y = "\pm" "\sqrt{3}x"


hence,

Df(x,y) is not invertble at (a,0) "\in" R2 for any a "\in" R also Df(x,y) is not invertble at (x, "\pm\sqrt{3}x") "\in" R2

"\implies" f(x,y) is not invertible.


at point (2,1)

Df(2,1) = 2(1)2[(1)2 "-" 3(2)2]

Df(2,1) = 2(1 "-" 12)

Df(2,1) = -22 "\neq" 0

hence, f(x,y) is locally invertible at (2,1)